OpenStudy (anonymous):

14. What is the slope of the line that passes through the points (3, 1) and (8, 3)?

OpenStudy (texaschic101):

slope = (y2 - y1) / (x2 - x1) (3,1)...x1 = 3 and y1 = 1 (8,3)...x2 = 8 and y2 = 3 now we sub slope = (3 - 1) / (8 - 3) slope = 2/5 any questions ?